OpenStudy (mmend98):

A metal sphere has a charge of -6 µC. How many excess electrons are there on the sphere?

OpenStudy (mmend98):

@TheSmartOne

OpenStudy (mmend98):

@Luigi0210 thanks guys in advance . this is the only question I have

OpenStudy (sharmagk83):

Here q = -6 uC, Charge on an electron is \[e = -1.6\times 10^{-19}C\] We know that \[q = n e\] \[n = \frac{ q }{ e } = \frac{ -6\times10^{-6}C }{ -1.6\times 10^{-19} C} = 3.75\times 10^{13}\]

OpenStudy (osprey):

yup, sharmagk83 is right. dividing the total charge by the basic building block of charge - the charge on the electron or the proton gives the number of charges in the sample. This could be said to show that charge is quantised - with thoughts of quantum theory looming. The same trick can be pulled for the idea of magnetisation, where this time the basic building block of magnetisation is the Bohr magneton. Named after Niels Bohr, Danish Nobel Prize winner for his theory of the the structure of the atom.